OpenStudy (anonymous):

A plane is headed N70E at an airspeed of 1000kph. From the ground it is determined that the resultant velocity of the place is 1010kph in the direction of N65E. Use Cartesian vectors to find the velocity of the wind.

OpenStudy (anonymous):

N70E - does that mean 70º north of east?

OpenStudy (anonymous):

no its 70 east of north

OpenStudy (anonymous):



OpenStudy (anonymous):

v_r = relative to ground v_a = actual the line that completes the triangle is velocity of wind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i do from here?

OpenStudy (anonymous):

by triangle law of vector addition, we have v_r=v_a+v_w do for x and y components independently

OpenStudy (anonymous):

i mean, v_a=v_r+v_w which gives us: v_w = v_a-v_r v_ax = 1010*cos(90-65), v_ay=1010*sin(90-65) v_rx = 1000*cos(90-70), v_ry = 1000*sin(90-70) v_wx = -65 v_wy = 162.98

OpenStudy (anonymous):

my bad.. v_wx= -24.32 v_wy = 84.82

OpenStudy (anonymous):

v_w = 88.24 kmph @ 74° South of E

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

sorry i was away

OpenStudy (anonymous):

does v_r mean v dot r?

OpenStudy (anonymous):

oh never mind

OpenStudy (anonymous):

yes thankyou for the help!